OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to give line express in y=mx+b (5,9), x+7y=9

OpenStudy (saifoo.khan):

parallel means same slope.

OpenStudy (saifoo.khan):

meme, wanna try?

OpenStudy (mimi_x3):

lols , idk i might get it wrong bcos i suck at these type of questions

OpenStudy (anonymous):

should i?

OpenStudy (anonymous):

fine saifoo u do it

OpenStudy (saifoo.khan):

c'mon u do it.

OpenStudy (anonymous):

y2-y1=m(x2-x1)

OpenStudy (anonymous):

y-9=7(x-5) y=7x-35+9

OpenStudy (anonymous):

y=7x-26

OpenStudy (anonymous):

or write it as 7x-y=26

jimthompson5910 (jim_thompson5910):

slope isn't 7 hahd, it's -1/7

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

see thats why i told saifoo to do it, im extremly sleepy

jimthompson5910 (jim_thompson5910):

Alternative route: Any equation parallel to x+7y=9 is of the form x+7y=C x+7y=C 5+7(9)=C 5+63=C 68=C C=68 So the parallel equation is x+7y=68

jimthompson5910 (jim_thompson5910):

now solve for y to get \[\large y = -\frac{1}{7}x+\frac{68}{7}\]